# A Constructive Definition of Space via Dynamical Evolution and Observational Acts

### v1

@@ 1 Introduction
A related line of inquiry can be found in Connes’ noncommutative geometry [5], which extends the Gel’fand–Naimark duality—originally linking commutative $C^{*}$-algebras and topological spaces—to the noncommutative setting. In Connes’ framework, geometry is encoded algebraically via $C^{*}$-algebras, leading to the notion of “quantum spaces.” However, this formulation lacks an explicit account of the dynamical and observational aspects that are central to our approach. In contrast, we argue that it is precisely the time evolution of a system and the observer’s choice of observable that jointly give rise to spatial structures. The topology of space is dynamically determined by the system’s evolution, while the measure structure emerges from the act of observation. In quantum theory, this leads naturally to the context-dependence of mutually incompatible observables—i.e., the impossibility of simultaneous measurement. While Connes’ formalism assumes a fixed topological and measure-theoretic backdrop within a $C^{*}$-dynamical system, our proposal allows these structures to vary depending on the dynamical system and the observer’s choices.
@@ 2.1 Dynamical Systems
Here, $U$ is a function representing the dynamical evolution. Although this notation may resemble the unitary evolution operator in quantum theory, it should be noted that $U$ is not assumed to be unitary in general. That is, we are not restricting ourselves to Hamiltonian systems.
@@
Although one could consider more exotic scenarios—such as systems with multiple time parameters or higher-dimensional time—we restrict ourselves here to the simplest case: a single, linear notion of time.
@@
If time is modeled instead by a group, the existence of inverse elements allows for time-reversal symmetry: for every $t$, there exists $-t$ such that $U(-t)$ is inverse of $U(t)$.
@@ 2.2 Reachability Domains
$\bar{X}_{0}^{T}(U)\equiv D(X_{0},T)(U)\equiv\{D(x_{0},T)|x_{0}\in X\}$ (2.7)
=$=\{U(x_{0},t)|x_{0}\in X,0\leq t\leq T\}.$ (2.8)
We use the overline notation $\bar{X}_{0}^{T}$ to suggest a closure operation, anticipating the later identification of this object with a kind of dynamical closure in the context of pre-topological structures.
@@ 2.4 Dynamical Interpretation of Pre-Topology
A pre-topological space on a set $X$, denoted as ($(\{X_{0}\subset X\},U,T)$, is defined by equipping $X$with a dynamical system $U$ and a time horizon $T$, and by examining how subsets $X_{0}\subset X$ evolve over time.
@@ 3.1 Measurable Spaces
[…]
@@
A pair $(X,\mathcal{B}$), where $\mathcal{B}$ is $\sigma$-algebra of subsets of $X$, is called a measurable space. In our context, this $\sigma$-algebra reflects the observer’s selection of which distinctions (i.e., which properties) are relevant and observable.
@@ 3.3 Introducing Measures: The Foundation of the Concept of State
The expectation value of an observable $f(\hat{x})$ , given a measure $\mu$, is written as:
[…]
@@ 3.4 Noncommutative Space and Quantum Theory
When an observer selects a particular observable $\hat{x}$, only those observables that commute with $\hat{x}$ can meaningfully define a simultaneous measurement context. The corresponding measure space $(X(\hat{x}),\mathcal{B}(\hat{x}),\mu(\hat{x}))$ emerges as the projection of the system onto this selected observable context. In other words, the “space” that becomes visible to the observer is the one constructed by the act of choosing $\hat{x}$.
@@ 3.5 Constructive Dynamics and Measurement in Spin Precession
As a concrete example, we consider the precession of a spin in an external magnetic field. We demonstrate how the dynamical evolution of the spin gives rise to a reachability domain, and how the observer’s choice of observable affects not only the associated measurable structure but also the reachability domain itself.
@@
$[\sigma_{i},\ \sigma_{j}]=2i\epsilon_{ijk}\sigma_{k},\ \{\sigma_{i},\sigma_{j}\}=2\delta_{ij}.$ (3.17)
@@
$[\hat{\psi}_{i},\ \hat{\psi}_{j}]=i\hbar\epsilon_{ijk}\hat{\psi}_{k},\ \{\hat{\psi}_{i},\hat{\psi}_{j}\}=\frac{\hbar^{2}}{2}\delta_{ij}.$
@@
This result means the following: if the observable is co-rotated with the state—that is, we choose $\hat{\psi}_{3}(t)$ as the observable at time $t$—then the system remains in an eigenstate with eigenvalue /2Planck-constant-over-$-\hbar/2$ at all times. In other words, by rotating the observable $\hat{\psi}_{3}$ in synchrony with the precession of the spin (in fact, at twice the angular speed), one can continuously extract the same measurement result without disrupting the reachability domain.
@@ Appendix A The GNS Construction and the Constructive Understanding of Space
where $|\hat{y}\rangle$ and $\langle\hat{x}|$ are understood as elements of a vector space and its dual, respectively.
@@ Appendix B Commutativity and the Emergence of Classical Space: A Reconsideration via Quantum Observation
$\omega(\hat{x})=\langle\Omega_{\omega}|\pi_{\omega}(\hat{x})|\Omega_{\omega}\rangle=\mathrm{Tr}|\Omega_{\omega}\rangle\langle\Omega_{\omega}|\pi_{\omega}(\hat{x})\equiv\mathrm{Tr}\left(\hat{\rho}_{\omega}\pi_{\omega}(\hat{x})\right).$ (B.2)
@@
$=\sum_{x\in B(\hat{x})}x\rho(x,x)dx\equiv\sum_{x\in B(\hat{x})}xd\mu(\hat{x}).$ (B.4)
Here, $\mathcal{B}(\hat{x})$ is $\sigma$-algebra generated by the spectral decomposition of $\hat{x}$—that is, by the set of its eigenvalues and corresponding eigenspaces. Let us denote the underlying “space” on which this $\sigma$-algebra is defined as $X(\hat{x})$.

### v2

@@ 1 Introduction
A related line of inquiry can be found in Connes’ noncommutative geometry [5], which extends the Gel’fand–Naimark duality—originally linking commutative $C^{*}$-algebras and topological spaces—to the noncommutative setting. In Connes’ framework, geometry is encoded algebraically via $C^{*}$-algebras, leading to the notion of “quantum spaces.” However, this formulation lacks an explicit account of the dynamical and observational aspects that are central to our approach. In contrast, we argue that it is precisely the time evolution of a system and the observer’s choice of observable that jointly give rise to spatial structures. The topology of space is dynamically determined by the system’s evolution, while the measure structure emerges from the act of observation. In quantum theory, this leads naturally to the context-dependence of mutually incompatible observables—i.e., the impossibility of simultaneous measurement. While Connes’ formalism assumes a fixed topological and measure-theoretic backdrop within a $C^{*}$-dynamical system, our proposal allows these structures to vary depending on the dynamical system and the observer’s choices [^1] While the essential contribution of Connes’ approach is the introduction of the Dirac operator to define geometric structures, his framework does not explicitly address the dynamical and observation-dependent aspects emphasized here, and is thus more static in character..
The groundbreaking aspect of Connes’approach lies in the observation that, once the algebra of “functions” on space becomes noncommutative, the very notion of a point loses its conventional meaning. In this sense, noncommutative geometry deals with what is often called a point-less geometry. By contrast, in our constructive framework, the set of observables under consideration is always taken to be a mutually commuting set within a chosen context, so that within this restricted set the assignment of values is still meaningful. In this sense, the notion of a “point” retains a partial operational meaning.
@@
As another line of discussion, following Born’s proposal of the reciprocity principle[14], there has been recent interest in quantum gravity scenarios that take seriously the possibility that momentum space itself may be curved[15][16]. Such approaches often regard phase space, rather than spacetime, as the fundamental arena, with the dynamical nature of the observer influencing the reconstruction of space[17]. This perspective shares common ground with our view that space should be constructed from dynamics and observation. However, from our standpoint, momentum is regarded merely as one possible observable, and in this sense our formulation may be considered more general.
@@ 2.1 Dynamical Systems
Here, $U$ is a function representing the dynamical evolution. Although this notation may resemble the unitary evolution operator in quantum theory, it should be noted that $U$ is not assumed to be unitary in general. That is, we are not restricting ourselves to Hamiltonian systems [^3] Note that the only primitives assumed here are a set and its dynamics; no metric or pre-existing topological structure is presupposed.
@@
Although one could consider more exotic scenarios—such as systems with multiple time parameters or higher-dimensional time—we restrict ourselves here to the simplest case: a single, linear notion of time444 Generalization to the case where time is modeled as a poset is technically straightforward, and would naturally connect to Lieb-Robinson-type causal structures in many-body quantum systems, suggesting further links to Lieb’s thermodynamic formulation[18]. A detailed treatment is beyond the scope of the present work..
@@
If time is modeled instead by a group, the existence of inverse elements allows for time-reversal symmetry: for every $t$, there exists t-t such that U(t)U(-t) is inverse of $U(t)$.
@@ 2.2 Reachability Domains
$\bar{X}_{0}^{T}(U)\equiv D_{U}(X_{0},T)\equiv\{D_{U}(x_{0},T)|x_{0}\in X_{0}\}$ (2.7)
$=\{U(x_{0},t)|x_{0}\in X_{0},0\leq t\leq T\}=\mathrm{cl}_{U}^{T}(X_{0}).$ (2.8)
Here,
$D_{U}(x_{0},T):=\{\,U(x_{0},t)\mid 0\leq t\leq T\,\}.$ (2.9)
We use the overline notation $\bar{X}_{0}^{T}$ as a shorthand for $\mathrm{cl}_{U}^{T}(X_{0})$, to suggest a closure operation, anticipating the later identification of this object with a kind of dynamical closure in the context of pre-topological structures. Here $\mathrm{cl}_{U}^{T}$ plays the role of a closure operator; if it satisfies idempotency, the induced structure is a topology, otherwise a pre-topology.
@@ 2.4 Dynamical Interpretation of Pre-Topology
A pre-topological space on a set $X$, denoted as $(\{X_{0}\subset X\},U,T)$, is defined by equipping XXwith a dynamical system $U$ and a time horizon $T$, and by examining how subsets $X_{0}\subset X$ evolve over time.
@@ 3.1 Measurable Spaces
where $\not${O} is the empty set. These subsets encode all the possible distinctions that can be made by combining the two properties.
@@
A pair $(X,\mathcal{B}$), where $\mathcal{B}$ is $\sigma$-algebra of subsets of $X$, is called a measurable space [^5] The point here is not the textbook definition of a $\sigma$-algebra per se, but the interpretation of this structure as arising from distinctions instituted through acts of observation.. In our context, this $\sigma$-algebra reflects the observer’s selection of which distinctions (i.e., which properties) are relevant and observable.
@@ 3.3 Introducing Measures: The Foundation of the Concept of State
The expectation value of an observable f(x^)f($\hat${x}) , given a measure $\mu$, is written as:
$\omega(f(\hat{x}))\equiv\langle f(\hat{x})\rangle_{\rho}=\int f(x)d\mu(x)=\int f(x)\rho(x)dx$ (3.13)
@@ 3.4 Noncommutative Space and Quantum Theory
When an observer selects a particular observable $\hat{x}$, only those observables that commute with $\hat{x}$ can meaningfully define a simultaneous measurement context. The corresponding measure space $)(X(\hat{x}),\mathcal{B}(\hat{x}),\mu(\hat{x}$)) emerges as the projection of the system onto this selected observable context. In other words, the “space” that becomes visible to the observer is the one constructed by the act of choosing $\hat{x}$.
@@ 3.5 Constructive Dynamics and Measurement in Spin Precession
As a concrete example, we consider the precession of a spin in an external magnetic field. We demonstrate how the dynamical evolution of the spin gives rise to a reachability domain, and how the observer’s choice of observable affects not only the associated measurable structure but also the reachability domain itself [^6] While the analysis of spin precession is standard, our discussion differs in that it interprets the resulting trajectories as reachability domains whose structure depends on the choice of observable, thereby providing a concrete example of space emerging from dynamics and measurement..
@@
$\sigma_{i},\ \sigma_{j}]=2i\epsilon_{ijk}\sigma_{k},\ \{\sigma_{i},\sigma_{j}\}=2\delta_{ij}.$ (3.17)
@@
$\hat{\psi}_{i},\ \hat{\psi}_{j}]=i\hbar\epsilon_{ijk}\hat{\psi}_{k},\ \{\hat{\psi}_{i},\hat{\psi}_{j}\}=\frac{\hbar^{2}}{2}\delta_{ij}.$
@@
This result means the following: if the observable is co-rotated with the state—that is, we choose $\hat{\psi}_{3}(t)$ as the observable at time tt—then the system remains in an eigenstate with eigenvalue $\hbar/2$ at all times. In other words, by rotating the observable $\hat{\psi}_{3}$ in synchrony with the precession of the spin (in fact, at twice the angular speed), one can continuously extract the same measurement result without disrupting the reachability domain.
@@ Appendix A The GNS Construction and the Constructive Understanding of Space
[…]
@@ Appendix B Commutativity and the Emergence of Classical Space: A Reconsideration via Quantum Observation
[…]
@@
[…]